OpenStudy (anonymous):

If x is a number that is 7 more than y, how do you express x as a function of y?

OpenStudy (anonymous):

a. x = f(y) = 7 + x b. x = f(y) = 7 + y c. y = f(x) = 7x d. y = f(x) = y – 7 e. y = f(x) = 7 + y